girls test scores in 1998 averaged 159, boys were 140, by 2002 girls were 160, boys were 136 so in 4 years the girls went up 1 pt. and the boys went down 4.Here is the formula given Girls W= 1/4x + 159. Boys W= -x + 140.W is the average writing score X years after 1998.
1.If current trends continue, when will the average test score for girls be 49 points higher than for boys?
2.If the current trends continue, when will the average test score for girls be 54 points higher than for boys?
They printed the answer for #1 as the year 2022, but I don't know how to get it. If I could answer the 1st then the second should be easy. I don't know how to set it up.

Answer by stanbon(75887)


4.Here is the formula given Girls W= 1/4x + 159. Boys W= -x + 140.W is the average writing score X years after 1998.
1.If current trends continue, when will the average test score for girls be 49 points higher than for boys?
2.If the current trends continue, when will the average test score for girls be 54 points higher than for boys?
They printed the answer for #1 as the year 2022 Girls W= (1/4)x + 159. Boys W= -x + 140.W is the average writing score X years after 1998.
1.If current trends continue, when will the average test score for girls be 49 points higher than for boys?
---
Equation:
Girls = Boys + 49
(1/4)x+159 = -x+140 + 49
(5/4)x = 30
(1/4)x = 6
x = 24
---
Year: 1998+24 = 2022
===========================
Cheers,
Stan H.
===========================